OpenStudy (anonymous):

The initial value of a car is $35,000. After a year, the value of the car is $31,850. Write an exponential function to model the expected value of the car. Estimate the value of the car after 5 years

OpenStudy (anonymous):

y(t) = c * a^t, t being time in years y(0) = c* a^0 = c*1 = c = 35 000 y(t) = 35000 * a^t y(1) = 31850 31850 = 35000 * a^1 31850 / 35000 = a = 0.91 a = 0.91 means that its value decreases by 1-0.91 = 9% each year y(5) = 35000 * 0.91^5 ~ 21 841

OpenStudy (anonymous):

is that the answer breh? and thanks

OpenStudy (anonymous):

$21 841 is the answer, yes :) If anything was unclear please say so.